the sum of the second and the fifth terms of an arithmetic sequence is 47,while the sum of the first eight terms is 116.Find the first two terms.thank you
Answer by richwmiller(17219)

2a + 5d=47,
116 = ½(2a + d(7))8
a = 46, d = -9
first two terms 46 and 37
2a + 5d=47,
116 = 4(2a + 7d)
116=8a+28d
2a + 5d=47
S = ½(92 -63)8
S=4*(29)
46,37,28,19, 10, 1, -8, 17
add to 116